OpenStudy (anonymous):

7. Solve the following trinomial: + 9x + 18 = 0

OpenStudy (mimi_x3):

(x+3)(x+6) x = -3 , -6

OpenStudy (anonymous):

\[x^2+9x+18=0\] \[x^2+3x+6x+18=0\] \[x(x+3)+6(x+3)=0\] \[(x+3)(x+6)=0\] either x+3=0 or x+6=0 hence \[x+3=0,x=-3\] \[x+6=0,x=-6\]